Tim Trepid is highly risk-averse while Mike Macho actually enjoys taking a risk.

a. Which one of the four investments should Tim choose? Compute coefficients of variations to help you in your choice.

a. Which one of the four investments should Tim choose? Compute coefficients of variations to help you in your choice.

Investments                            Returns:                                  Standard

                                            Expected Value                            Deviation

Buy Stocks.............................$9,140                                    $6,140

Buy bonds................................7,680                                       2,560

Buy commodity future............19,100                                      26,700

Buy options.............................17,700                                     18,200

b. Which one of the four investments should Mike choose?
Answer #1

Standard Deviation oeffi variation = cient of Expected Return

Buy Stocks:

Coefficient of variation = 6140/9140

=0.6712

Buy bonds:

Coefficient of variation = 2560/7680

=0.333

Buy commodity future:

Coefficient of variation = 26700/19100

=1.398

Buy options:

Coefficient of variation = 18200/17700

1.028

Hence, Tim should purchase bonds because the lower the ratio of standard deviation to mean return, the better your risk-return tradeoff